# Center Biophysics and Biochemistry, (IVIC), Caracas, Venezuela

### v2

@@ Abstract
Complexity is a fundamental characteristic of Our (The?) Universe. Without complexity life will not exist [10].
@@
1.2 Complexity, uncertainty and esperimental error.
@@
A common view is to consider data dispersion as ‘errors’, something related with experimenters or equipment ‘mistakes’, which certainly do occur sometimes. Yet, an extremely important source of data uncertainty, is not ‘error’, but a property of the system under study. Such is the case of quantum physics [94] and related disciplines. But fuzziness of data is fundamental importance in biology, if all elements of any population are equal, they can by exterminated by noxae such as a pandemic, or by defects resulting from inbreeding in a small uniform population. When a species is reduced to small sets of individuals, the species in condemned to extinction. Many fundamental biological signals (hearth beat, brain signals) in healthy individuals are fuzzy and chaotic, and brcome less chaotic in pathological states like heart beat [99, 100, 145] or electroencephalographic (EEG) signals [Shamsi2021, Spacic2011, 115, 38, Klonowski2005, 128, Spasic2011a, 90, 49, Perez2022, 87] (a non exhaustive list of examples). Fractal dimension is also used to analyze lung sounds [72, 71, 88]. Besides biomedical uncertainty, there is a large number of papers using fractal dimension in subjects such as geomagnetic field studies [76], mammary [162], ultrasound studies [45] and machines and materials failures [82, ValtierraRodriguez2019] and other situations [59, Sharma2013, RodriguezHernandez2022]. Despite the abundance of publications where fractal dimension is calculated, the authors often do not try to understand the mathematics.
In this review we will consider several proposed modes to calculate the fractal dimension [96, 97, 114, Sevcik1998a, Sevcik2010], and to present several examples of ‘dispersion’ related to chaotic nonlinear systems which are some sources of fractality. We will also consider the Hurst´s coefficient [105] for which a simple relationship whih $D$ was proposed [137, 92], a relation currently considered wrong [141, 68, 69, Sutcliffe2016].
@@
Complex systems are often called nonlinear systems, strongly sensitive to initial conditions, systems that are also known as ‘dynamical systems’. In these there is no randomness, but they change as a result of the accuracy of the calculations or very small environmental variations. An example is the climate [131, 130, 159], where very small changes can produce very large effects, the flutter of a butterfly in China can produce a storm in America, the so-called butterfly effect, poetic name that is reinforced by the shape of the Lorenz attractor [131] (See Figure 1). Dynamical systems are totally deterministic, not random, but they are unpredictable [131, 37, 129, 110].
Another form of uncertainty is called chaos. It is central to all fields of human knowledge including quantum physics [26, Zurek20003, 101]. This chapter is an introduction to chaos and its difference from the statistical uncertainty to which the rest of this book relates. It is not a new concept started with Lorenz almost 60 years ago [131] and remains central in almost all areas of scientific knowledge.
The impact of Lorenz ants butterfly effect on weather prediction, has determined that the concept of chaos is usally said that vas created by Lorenz [131], still the concept is earlier, it was introduced by Turing [Turing1952] to explain the source of biological structure and complexity in biology, and by Belousov [24, 104, Zhabotinsky2007].
Another result of chaos in nonlinear systems is turbulence [161, 113]. A thin layer of ice on an airplane’s wings can cause enough turbulence to prevent it from flying. Chaos exists in vital functions such as the normal heart rate which is, within certain limits, chaotic, the absence of chaos in the heart rate indicates disease [163, 73, 109, 166, 164, 99, 165], but if the chaoticity becomes extreme (the so-called ventricular fibrillation) it causes death. Examples of the significance of chaos are too many to cite here, this includes earthquakes [52], fluctuations in the stock market and the economy in general [133, 133, 140, 123], a couple of somewhat classic references are [137, 92].
1.4 Lorenz´s Uncertainty.
But uncertainty is not just “error”. his was first shown in a log neglected model of weather studied by Lorenz [131] (Srr the system of Eqs. (1)). Lorenz´s system of equations was cwrtainly sensitive to rounding errors, and the limited accuracy of the analog computer available to him. But the strange set of solutions (resembling a butterfly), although constrained to a definite volume of an Euclidean space,never crossed it self at a previous point (solution). The system was also unpredictable, it changed if was initialized whith apparently similar values.
@@
All possible values in a dynamic system usually usually exist in a finite space and are distributed in a region of that space called attractor. A classical example id Lorenz´s attractor [131], shown in Figure 1 and represents the solutions of a simple set of differential equation built by Lorenz as a an atmospheric climate model. Lorenz´s equations are [131, 159, 130]:
@@
where x, y and z are the coordinates of an Euclidean system. Figure 1 represents a series of solutions of this system. The trace in the figure represents a single continuous, which never crosses through the same, previous, point, and repents unpredictable displacements of which are solutions of the equation´s system. Perhaps the most important properties of the graphic, is that all solutions are confined in a subset (volume) of the Euclidean space called a strange attractor, Equation system (1) was built by Lorenz as a climate model, is the first es el primer strange attractor known and to describe climate (we also call it weather) and it showed that real weather (much more complex than the model) is unpredictable (except for short periods), changing the history of meteorology. Having a strange attractor is not unique to Lorenz´s system, many other systems have strange attractors too.
Climate unpredictability has been well demonstrated after Lorenz, weather can only be predicted for short periods even with modern supercomputers of today, receiving information from sensors spread all over the world. Lorenz´s equations in laser models [89], electricity dynamos [116], convection loops [75],direct current motors without brushes [95],electric circuits [50] and chemical reactions[168] among many more systems. Lorenz´s -like systems are not this books´s object of study but they must be considered by its reader, since statistical methods are often applied to Lorenz-like dynamic systems, and are used to predict the properties of such systems, and the predictions failure is wrongly attributed to statistics and not to their unpredictable nature. An example of this opinion studies, which when are made public, may (usually d) modify the public opinion that they objectively claim to study. In politics, it sometimes results it is common ‘to bet on the winner’ voting for or against him or her if a candidate produces fear. Classical examples of this effect is also the expectation produced by a medical treatment (which may lead to changing or selecting the the subject of the study or the experimenters performing the study), they could even modify the results observed (i.e. change the result). increasing de beneficial or adverse effects of the treatment studied, having or not having faith in the treatment could be another example. This is an extension of the quantum physics observer effect, extended to daily supra molecular. Lorenzian uncertainty was later called chaos, and was the firs system known to contain chaotic uncertainty.
@@
We define here as time series what, perhaps in better but longer common English, should be called: series of events that occurs in time. In terms of a graph it would be a graph where the ordinate represents random events and the abscissa is the time at which each one occurs.
@@
The English term waveform, which we translate here as “waveform” refers to the appearance of a wave much more complex than a “ripple” as defined by the Meriam Webstwe´s Dictionary, usually periodic versus time. Any waveform is a series of points. Apart from classical statistical models such as statistical moments [85] and regression analysis, properties such as Kolmogorov-Sinai entropy entropy [79], the apparent enthronement [166] and the fractal dimension [114] have been proposed to address the problem of analyzing waveforms. The fractal dimension can provide information about spatial extension (tortuosity or its ability to fill space) and its self-similarity (ability to remain unchanged when the measurement scale changes), its self-affinity [20]. Unfortunately, although there are rigorous methods for calculating the fractal dimension [80, 81, 16, Ripoli1999], their usefulness is severely limited since they demand great computing power and because their evaluation is time-consuming. In Euclidean space, waveforms are flat, two-dimensional curves.
@@
According to Mandelbrot [137, pg. 15 and Chap. 39]:
@@
The term waveform applies to the form of a wave usually drawn as a value at an instant, of a periodic nature, versus time. A classic example is momentum statistics and regression analysis, properties such as the Kolmodorov-Sinai entropy entropy [79], the apparent entropy [166] and the fractal dimension [Sevcik1998a, Sevcik2010] have been proposed to perform waveform regression analysis. the fractal dimension can provide information on statistical extent (tortuosity or the ability to fill space) and similarity (the ability to remain unchanged when the measurement scale changes) and self-affinity [20]. In processes occurring in two Euclidean dimensions, waveforms are curves with coordinates, $x$ and $and$, which usually have different units.
@@
For a number of waveforms111$\in$ = belongs to; $\land$ = and; $\notin$ = does not bellong; $\mathbb{R}$ = set of all real numbers; $\mathbb{C}$ = set of all complex numvers.
@@
where $\mathbb{R}$ is the ser of real numbers, an $\mathbb{C}$ is the set of all real numbers. which does not necessarily mean that $D_{HB}>1$. There are also functions such as, for example, the call ‘dust’ or Cantor set [39, Darst1993, 54, 8, 11] for which $D_{HB}=\frac{log(2)}{log(3)}=0.630929753571\ldots$. These waveforms correspond to infinite disjoint sets of points which, however, constitute a waveform [11].
@@
where $d$ is the planar extension of the [137, Chapter 12] curve and $L$ is the length of the discretized curve defined as:
@@
3.2 A simple method to calculate the fractal dimension of waveforms Sevcik´s [Sevcik1998a, Sevcik2010] fractal dimension.
An expression to calculate the fractal dimension of a waveform is obtained from the fractal dimension of Hausdorff–Besicovitch ($D_{HB}$) [93, 27]. The definition of fractal of Mandelbrot (see for example [138]) to see that the Hausdorff-Besicovitch dimension is not an integer. The Hausdorff-Besicovitch dimension [138] of a metric space (for a very understandable discussion of metric spaces see Barnsley [20]) can be expressed as:
@@
where the notation is equal to that of Eq. (2). In a metric space given any point $P$, an open ball with center at $P$, $\epsilon$, is the set of all points $x$ for which $dist(P,x)<\epsilon$, at any length $L$ can be divided into $N(\epsilon)=L/(2\cdot\epsilon)$ long segments ${2\cdot\epsilon}$, and any of them be covered by $N$ open balls of radius $\epsilon$. Therefore the Eq. (6) can be rewritten as
$\begin{split}D_{HB}&=\underset{\epsilon\rightarrow 0}{\lim}\left[\frac{-\ln(L)+\ln(2\cdot\epsilon)}{\ln(\epsilon)}\right]\\ \cdots&=\underset{\epsilon\rightarrow 0}{\lim}\left[1-\frac{\ln(L)-\ln(2)}{\ln(\epsilon)}\right]\\ \boldsymbol{{\text{\wasytherefore{\quad}}D_{HB}}}&=\boldsymbol{{\underset{\epsilon\rightarrow 0}{\lim}\left[1-\frac{\ln(L)}{\ln(\epsilon)}\right]\quad\boldsymbol{\square}}}\end{split}$ (7)
Waveforms are flat caves eb ub space with coordinates with distinct units. Since the topology of a metric space does not change under linear transformations, it is convenient to linearly transform one waveform into another into a normalized space, where all axes are equal. This can be done with two and linear transformations in another embedded in an equivalent metric space. The first transformation normalizes the curve in the abscissa as:
@@
the approximation to $\Phi$ expressed in Eq. (7) expressed in Eq. (10), improves as $N^{\prime}\rightarrow\infty$. The Eq. (10) as simply $D$, in [Sevcik1998a, Sevcik2010] as an anonymous derivative of the Hausdorff-Besicovitch dimension [27, 93], this fractal dimension estimator has been called with increasing frequency: “Sevcik’s dimension”[Sharma2013, 53, Shi2018, 153, Xue2020, 119]. So since our work of 2022 [RodriguezHernandez2022] we call it “Sevcik´s fractal dimension”, $D_{S}$.
@@
Although $\Phi$ is a topological invariant of a set and a metric space, $D_{S}$ is only an empirical estimate of $\Phi$ with some uncertainty based on a set of points sampled from a wave; $D_{S}$ is therefore a random variable. The ratio between $\Phi$ and $D_{S}$ is similar to that between a mean of a population ($\mu$) and the mean $\bar{x}$ estimated when sampling the population; Although $\mu$ is a population invariant, $\bar{x}$ will change with sampling. Just as $\bar{x}$ converges towards $\mu$ as the sample approaches population size, $D_{S}$ converges towards $\Phi$ as $N^{\prime}\rightarrow\infty$. Now we will derive an expression for $var(D_{S})$ (the variance of the $D_{S}$) from the estimate of $D_{S}$ obtained by mastering $N^{\prime}$ points of a wave. It should be obvious from the derivation of $D_{S}$ and the non-stationary character of the values of $D_{S}$ determined with the Eq. (10), that ${\text{var}\left[{(}\right]}{D_{S}})$ does not provide information about the asymptotic value of D obtained as ${N^{\prime}\rightarrow\infty}$. The variance of $D_{S}$ can be estimated starting from the following expression:
$\begin{split}{\text{var}\left[{D_{S}}\right]}&={\text{var}\left[{\frac{\ln(L)}{\ln(2\cdot N^{\prime}))}}\right]}\\ \boldsymbol{{\text{\wasytherefore{\quad}}{\text{var}\left[{D_{S}}\right]}}}&=\boldsymbol{{{\text{var}\left[{\frac{\ln(L)}{\ln(2\cdot N\text{{\textquotesingle}})}}\right]}\quad\boldsymbol{\square}}}\end{split}$ (11)
The approximate solution to Eq. (11) can be obtained by recalling that the variance of any function of random variable sets $\{x_{i}\}$ approximates with a Taylor series (see for example [47])) as:
@@
3.2.2 Convergence of $D_{S}$ towards $D_{HB}$, an analytic solution for Koch’s triadic curve.
The asymptotic convergence of $D_{S}$ to $\Phi$ can be obtained from the Koch triad. Although I derived the equation (10) for waves, that is, plane curves that are sets of pairs of dots $(x_{i},y_{i})$ such as $x_{i}\rightarrow\infty$ when $i\rightarrow\infty$. However, on at least some occasions the utility of $D_{S}$ $\Phi$ extends to the field of waveforms, I will demonstrate that with the famous Koch triad shown in Figure 2.
Figure 2: Characterization of the Koch curve. The Koch curve is built on the three sides of an equilateral triangle, here we consider a single side to visualize better. We start with a line long 1, above and to the left of the graph (Stage 0 or s = 0, in the figure). Then at each stage the center of the segment removes the middle third of each lake line ${\lambda}$ and adds an equal segment of length ${\lambda}$/3 forming two sides of a triangle. Repeating this an infinite number of vece, one obtains Koch’s snowflake, each xstage is indicated by the letter $S$ increased by 1. Below and to the right of the figure is eta 3 ($S=2$) scalded into a square of side 1, in stage 3 ($S_{r}=3$), see the text for more details. Adapted from Figure 4 of [Sevcik1998a, Sevcik2010].
@@
$D_{H}=\underset{S\rightarrow\infty}{\lim}-{\frac{\ln\left(1+2^{{2S}}\right)}{\ln\left(\frac{3^{-S}}{2}\right)}}=\frac{\ln(4)}{\ln(3)}=1.2618\ldots$
@@
This is shown at the bottom of Figure 2 for the $3^{a}$ stage of the build process. The transformation does not modify the length of the horizontal components of the Koch curve but extends the length of all inclined sections that it becomes.
@@
as it should be. Thus, the lit of $\boldsymbol{{D_{S}}}$ is indeed $\boldsymbol{{\Phi}}$ when $\boldsymbol{{N\rightarrow\infty}}$ [Sevcik1998a, Sevcik2010].
3.2.3 Convergence towards $D_{HB}$ of Sevcik fractal dimension.
The main limitation of $D_{S}$ is that the speed of convergence towards $D_{HB}$ the latter is, generally speaking, unknown. Processes such as the one leading to a convergence such as that suggested by Eq. (17) suggest that convergence occurs, but says little or nothing about the speed of convergence.
Part of the problem is that the original convergence presumes that the set of pairs $\{x_{i},y_{i}\}$ that $\{x_{i}\}\in\mathbb{R}$ and that also $\{y_{i}\}\in\mathbb{R}$, but this is not necessarily always true. But what happens if we use the expressions (10) and (15) and $\{x_{i}\}\notin\mathbb{R}$ and that also ($(\{y_{i}\}\notin\mathbb{R})$?. This is the case when using the sdimension to determine whether the series of digits of $\pi$ is infinitely aperiodic, which in the field of aperiodic non-rational numbers is called normal (without, in this case, any relation to the Gauss pdf) [Sevcik2017b]:
@@
For any digit $c$ (on the scale of $g$) and each natural number $n\in\mathbb{N}$, we denote $I(c,n)$ the number of those digits rm the sequence $c_{1},c_{2},\ldots,c_{n}$, which are equal to $c$. Yes
$\overset{\lim}{n\rightarrow\infty}\left[\dfrac{I(c,n)}{n}\right]=\dfrac{1}{g}$
For each of the possible $9$ values of $c$, then the number $x$ called normal on the scale of $g$. A number that is natural on the $g$ scale is called absolutely natural [22]. For a number in base with $10$, the definition implies that $c_{i}\in\mathbb{N}[0,9]$ if it must be true for any number number $x\in\mathbb{R}$ such as $N\rightarrow\infty$. Therefore, various authors use statistical tests such as comparing the frequencies of each digit in the sequences of the decimals of $\pi$, and the frequencies of various combinations of decimal digits [19]. This approach is unsatisfactory since all possible combinations can be evaluated, and is limited by the $N$ values of other series studied, as well as by the size of subsamples of which $N$ was selected to perform statistical tests. This approach is unsatisfactory since not all possible combinations can be evaluated, and is limited by $N$ the values of the series studied, and by the size of the subsamples of each $N$ that is used to perform the statistical tests.
The fractal analysis described [Sevcik2017b] considered the set of decimals of $\pi$, and calculated its approximate fractal dimension for $N=10^{1},10^{2},10^{3},\ldots,10^{9}$ using the Sevcik´s fractal dimension (see specific details in [Sevcik2017b]) and showed that

### v3

@@ Abstract
Complexity is a fundamental characteristic of Our (The?) Universe. Without complexity life will not exist [HowManyUniverses2022].
@@
1.2 Complexity, uncertainty and experimental error.
@@
A common view is to consider data dispersion as ‘errors’, something related with experimenters or equipment ‘mistakes’, which certainly do occur sometimes. Yet, an extremely important source of data uncertainty, is not ‘error’, but a property of the system under study. Such is the case of quantum physics [96] and related disciplines. But fuzziness of data is fundamental importance in biology, if all elements of any population are equal, they can by exterminated by noxae such as a pandemic, or by defects resulting from inbreeding in a small uniform population. When a species is reduced to small sets of individuals, the species in condemned to extinction. Many fundamental biological signals (hearth beat, brain signals) in healthy individuals are fuzzy and chaotic, and brcome less chaotic in pathological states like heart beat [101, 102, 147] or electroencephalographic (EEG) signals [Shamsi2021, Spacic2011, 117, 39, Klonowski2005, 130, Spasic2011a, 92, 50, Perez2022, 89] (a non exhaustive list of examples). Fractal dimension is also used to analyze lung sounds [74, 73, 90]. Besides biomedical uncertainty, there is a large number of papers using fractal dimension in subjects such as geomagnetic field studies [78], mammary [164], ultrasound studies [46] and machines and materials failures [84, ValtierraRodriguez2019] and other situations [60, Sharma2013, RodriguezHernandez2022]. Despite the abundance of publications where fractal dimension is calculated, the authors often do not try to understand the mathematics.
In this review we will consider several proposed modes to calculate the fractal dimension [98, 99, 116, Sevcik1998a, Sevcik2010], and to present several examples of ‘dispersion’ related to chaotic nonlinear systems which are some sources of fractality. We will also consider the Hurst´s coefficient [107] for which a simple relationship whih $D$ was proposed [139, 94], a relation currently considered wrong [143, 70, 71, Sutcliffe2016].
@@
Complex systems are often called nonlinear systems, strongly sensitive to initial conditions, systems that are also known as ‘dynamical systems’. In these SYSTEMS, there is no randomness, but they change as a result of the accuracy of the calculations or due to very small environmental variations., The classical example is earth climate shown by Loeenz [133, 132] to be unpredictable, sometimes even at short lapses of time [131, 62, 32].
Lorence [133, 161], proved that very small changes can produce very large whether changes, ‘the flutter of a butterfly in China can produce a storm in America’, the so-called butterfly effect, poetic name that is reinforced by the shape of the Lorenz attractor [133] (See Figure LABEL:F:Lorenz). Dynamical systems are totally deterministic, not random, but they are unpredictable [133, 38, 131, 112].
Another form of uncertainty is called chaos. It is central to all fields of human knowledge including quantum physics [26, Zurek20003, 103]. This chapter is an introduction to chaos and its difference from the statistical uncertainty to which the rest of this book relates. It is not a new concept started with Lorenz almost 60 years ago [133] and remains central in almost all areas of scientific knowledge.
The impact of Loren butterfly effect on wheather prediction, has determined that the concept of chaos is usually said that was created by Lorenz [133], still the concept is earlier, it was introduced by Turing [Turing1952] to explain the source of biological structure and complexity in biology, and by Belousov [24, 106, Zhabotinsky2007].
Another result of chaos in nonlinear systems is turbulence [163, 115]. A thin layer of ice on an airplaneLorenz’s wings can cause enough turbulence to prevent it from flying. Chaos exists in vital functions such as the normal heart rate which is, within certain limits, chaotic, the absence of chaos in the heart rate indicates disease [165, 75, 111, 168, 166, 101, 167], but if the chaoticity becomes extreme (the so-called ventricular fibrillation) it causes death. Examples of the significance of chaos are too many to cite here, this includes earthquakes [53], fluctuations in the stock market and the economy in general [135, 135, 142, 125], a couple of somewhat classic references are [139, 94].
1.4 Lorenz’s Uncertainty.
But uncertainty is not just “error”. his was first shown in a log neglected model of wheather studied by Lorenz [133] (S the system of Eqs. (1)). Lorenz´s system of equations was certainly sensitive to rounding errors, and the limited accuracy of the analog computer available to him. But the strange set of solutions (resembling a butterfly), although constrained to a definite volume of an Euclidean space,never crossed it self at a previous point (solution). The system was also unpredictable, it changed if was initialized with apparently similar values.
@@
All possible values in a dynamic system usually usually exist in a finite space and are distributed in a region of that space called attractor. A classical example id Lorenz´s attractor [133], shown in Figure LABEL:F:Lorenz and represents the solutions of a simple set of differential equation built by Lorenz as a an atmospheric climate model. Lorenz´s equations are [133, 161, 132]:
@@
where x, y and z are the coordinates of an Euclidean system. Figure LABEL:F:Lorenz represents a series of solutions of this system. The trace in the figure represents a single continuous, which never crosses through the same, previous, point, and repents unpredictable displacements of which are solutions of the equation´s system. Perhaps the most important properties of the graphic, is that all solutions are confined in a subset (volume) of the Euclidean space called a strange attractor, Equation system (1) was built by Lorenz as a climate model, is the first es el primer strange attractor known and to describe climate (we also call it whether) and it showed that real whether (much more complex than the model) is unpredictable (except for short periods), changing the history of meteorology. Having a strange attractor is not unique to Lorenz´s system, many other systems have strange attractors too.
Climate unpredictability has been well demonstrated after Lorenz. Wheather can only be predicted for short periods even with modern supercomputers of today, receiving information from sensors spread all over the world. Lorenz´s equations in laser models [91], electricity dynamos [118], convection loops [77],direct current motors without brushes [97], electric circuits [51] and chemical reactions [170] among many more systems. Lorenz’s-like systems are not this books´s object of study but they must be considered by its reader, since statistical methods are often applied to Lorenz-like dynamic systems, and are used to predict the properties of such systems, and the predictions failure is wrongly attributed to statistics and not to their unpredictable nature. An example of this opinion studies, which when are made public, may (usually d) modify the public opinion that they objectively claim to study. In politics, it sometimes results it is common ‘to bet on the winner’ voting for or against him or her if a candidate produces fear. Classical examples of this effect is also the expectation produced by a medical treatment (which may lead to changing or selecting theproved that study subject or the experimenters performing the study), they could even modify the results observed (i.e. change the result). increasing de beneficial or adverse effects of the treatment studied, having or not having faith in the treatment could be another example. This is an extension of the quantum physics observer effect, extended to daily supra molecular. Lorenzian uncertainty was later called chaos, and was the firs system known to contain chaotic uncertainty.
@@
We define here as time series what, perhaps in better but longer common English, should be called: series of events which occur in time. A set presented as a graph where the ordinate represents random events and the abscissa is the time at which each event occurs.
@@
The English term waveform, which we translate here as “waveform” refers to the appearance of a wave much more complex than a “ripple” as defined by the Meriam Webstwe´s Dictionary, usually periodic versus time. Any waveform is a series of points. Apart from classical statistical models such as statistical moments [87] and regression analysis, properties such as Kolmogorov-Sinai entropy [81], the apparent entropy [168] and the fractal dimension [116] have been proposed to address the problem of analyzing waveforms. The fractal dimension can provide information about spatial extension (tortuosity or its ability to fill space) and its self-similarity (ability to remain unchanged when the measurement scale changes), its self-affinity [20]. Unfortunately, although there are rigorous methods for calculating the fractal dimension [82, 83, 16, Ripoli1999], their usefulness is severely limited since they demand great computing power and because their evaluation is time-consuming. In Euclidean space, waveforms are flat, two-dimensional curves.
@@
According to Mandelbrot [139, pg. 15 and Chap. 39]:
@@
The term waveform applies to the form of a wave usually drawn as a value at an instant, of a periodic nature, versus time. A classic example is momentum statistics and regression analysis, properties such as the Kolmodorov-Sinai entropy [81], the apparent entropy [168] and the fractal dimension [Sevcik1998a, Sevcik2010] have been proposed to perform waveform regression analysis. the fractal dimension can provide information on statistical extent (tortuosity or the ability to fill space) and similarity (the ability to remain unchanged when the measurement scale changes) and self-affinity [20]. In processes occurring in two Euclidean dimensions, waveforms are curves with coordinates, $x$ and $y$, which usually have different units.
@@
For a number of waveforms
@@
where $\mathbb{R}$ is the ser of real numbers, an $\mathbb{C}$ is the set of all real numbers. which does not necessarily mean that $D_{HB}>1$. There are also functions such as, for example, the call ‘dust’ or Cantor set [40, Darst1993, 55, 7, 9] for which $D_{HB}=\frac{log(2)}{log(3)}=0.630929753571\ldots$. These waveforms correspond to infinite disjoint sets of points which, however, constitute a waveform [9].
@@
where $d$ is the planar extension of the [139, Chapter 12] curve and $L$ is the length of the discretized curve defined as:
@@
3.2 Sevcik´s fractal dimension [Sevcik1998a, Sevcik2010]: a simple method to calculate the fractal dimension of waveforms.
An expression to calculate the fractal dimension of a waveform is obtained from the Hausdorff–Besicovitch ($D_{HB}$) dimension [95, 27]. Mandelbrot´s fractal (see for example [140]) is real number, not an integer. The Hausdorff-Besicovitch dimension [140] of a metric space can be expressed (for a very understandable discussion of metric spaces see Barnsley [20]) as:
@@
where the notation is equal to that of Eq. (2). In a metric space given any point $P$, an open ball with center at $P$, $\epsilon$, is the set of all points $x$ for which $dist(P,x)<\epsilon$, at any length $L$ can be divided into $N(\epsilon)=L/(2\cdot\epsilon)$ long segments ${2\cdot\epsilon}$, and any of them be covered by $N$ open balls of radius $\epsilon$. Therefore Eq. (6) can be rewritten as
$\begin{split}D_{HB}&=\underset{\epsilon\rightarrow 0}{\lim}\left[\frac{-\ln(L)+\ln(2\cdot\epsilon)}{\ln(\epsilon)}\right]\\ \cdots&=\underset{\epsilon\rightarrow 0}{\lim}\left[1-\frac{\ln(L)-\ln(2)}{\ln(\epsilon)}\right]\\ \boldsymbol{{\text{\wasytherefore{\quad}}D_{HB}}}&=\boldsymbol{{\underset{\epsilon\rightarrow 0}{\lim}\left[1-\frac{\ln(L)}{\ln(\epsilon)}\right]\quad\boldsymbol{\square}}}\end{split}.$ (7)
Waveforms in two diminutions are sets of points with coordinates which may have distinct units. Since the topology of a metric space does not change under linear transformations, it is convenient to linearly transform one waveform into another in a normalized space, where all axes are equal and have legth $1$. This can be done with two linear transformations embedded into an equivalent metric space. The first transformation normalizes the curve´s abscissa as:
@@
the approximation to $\Phi$ expressed in Eq. (7) expressed in Eq. (10), improves as $N^{\prime}\rightarrow\infty$. The Eq. (10) as simply $D$, in [Sevcik1998a, Sevcik2010] as an anonymous derivative of the Hausdorff-Besicovitch dimension [27, 95], this fractal dimension estimator has been called with increasing frequency: “Sevcik´s ” dimension [Sharma2013, 54, Shi2018, 155, Xue2020, 121]. So since our work of 2022 [RodriguezHernandez2022] we call it “Sevcik´s fractal dimension”, $D_{S}$.
@@
Although $\Phi$ is a topological invariant of a set and a metric space, $D_{S}$ is only an empirical estimate of $\Phi$ with some uncertainty based on a set of points sampled from a wave; $D_{S}$ is therefore a random variable. The ratio between $\Phi$ and $D_{S}$ is similar to that between a mean of a population ($\mu$) and the mean $\bar{x}$ estimated when sampling the population; Although $\mu$ is a population invariant, $\bar{x}$ will change with sampling. Just as $\bar{x}$ converges towards $\mu$ as the sample approaches population size, $D_{S}$ converges towards $\Phi$ as $N^{\prime}\rightarrow\infty$. Now we will derive an expression for $var(D_{S})$ (the variance of the $D_{S}$) from the estimate of $D_{S}$ obtained by mastering $N^{\prime}$ points of a wave. It should be obvious from the derivation of $D_{S}$ and the non-stationary character of the values of $D_{S}$ determined with Eq. (10), ${\text{var}\left[{(}\right]}{D_{S}})$ does not provide information about the asymptotic value of $D_{S}$ obtained as ${N^{\prime}\rightarrow\infty}$. The variance of $D_{S}$ can be estimated starting from the following expression:
$\begin{split}{\text{var}\left[{D_{S}}\right]}&={\text{var}\left[{1+\frac{\ln(L)}{\ln(2\cdot N^{\prime}))}}\right]}\\ \boldsymbol{{\text{\wasytherefore{\quad}}{\text{var}\left[{D_{S}}\right]}}}&=\boldsymbol{{{\text{var}\left[{\frac{\ln(L)}{\ln(2\cdot N\text{{\textquotesingle}})}}\right]}\quad\boldsymbol{\square}}}\end{split}$ (11)
The approximate solution to Eq. (11) can be obtained by recalling that the variance any random variable set may be $\{x_{i}\}$ approximated with a Taylor series (see for example [48])) as:
@@
3.2.2 Convergence of $D_{S}$ towards $D_{HB}$, an analytic solution for Koch´s triadic curve.
The asymptotic convergence of $D_{S}$ to $\Phi$ can be proven from the Koch triad. Although I derived the equation (10) for waves, that is, plane curves that are sets of pairs of dots $(x_{i},y_{i})$ such as $x_{i}\rightarrow\infty$ when $i\rightarrow\infty$. However, on at least some occasions the utility of $D_{S}$ $\Phi$ extends to the field of waveforms, I will demonstrate that with the famous Koch triad shown in Figure 2.
Figure 2: Characterization of the Koch curve. The Koch curve is built on the three sides of an equilateral triangle, here I consider a single side to better visualize the convergence $D_{S}\rightarrow D_{HB}$ for the Koch´s snowflake. I start with a line of length $1$. Above and to the left of the graph (Stage 0 or $S=0$, in the figure) a line of legth 1 ($\lambda=1$) is presented. Then, at each stage, a segment of length $\frac{\lambda\cdot S}{3}$ is cut and moved up from each straight line, and an equal segment of length $\frac{\lambda\cdot S}{3}$ is added forming two sides of a new “triangle” lackin its “basal” segment. Repeating this an infinite number of times, one obtains Koch´s snowflake, each $S$ stage is indicated by the letter $S=0,1,2,\ldots,\infty$ Below and to the right of the figure, stage 3 ($S=2$) is shown scalded into a square of side $1$, in stage 3 ($S_{r}=3$), see the text for more details. This figure was adapted from Figure 4 of [Sevcik1998a, Sevcik2010].
@@
$D_{HB}=\underset{S\rightarrow\infty}{\lim}-{\frac{\ln\left(1+2^{{2S}}\right)}{\ln\left(\frac{3^{-S}}{2}\right)}}=\frac{\ln(4)}{\ln(3)}=1.2618\ldots$
@@
This is shown at the bottom of Figure 2 for the $3^{S}$ stage of the build process. The transformation does not modify the length of the horizontal components of the Koch curve but extends the length of all inclined sections that it becomes.
@@
as it should be. Thus, the limit of $\boldsymbol{{D_{S}}}$ is indeed $\boldsymbol{{\Phi}}$ when $\boldsymbol{{N\rightarrow\infty}}$ [Sevcik1998a, Sevcik2010].
3.2.3 Sevcik´s fractal dimension convergence to $D_{HB}$.
The main limitation of $D_{S}$ is that the speed of convergence towards $D_{HB}$, the latter is, generally speaking, unknown. Processes such as the one described by Eq. (17) suggest that convergence occurs, but say little about the speed of convergence.
Part of the problem is that the original convergence presumes that in the set of pairs $\{(x_{i},\in\mathbb{R})\;\boldsymbol{{\land}}\;(y_{i}\in\mathbb{R})\}$, but this is not necessarily always true. Then: What happens, if we use expressions (10) and (15) but $\{(x_{i},\notin\mathbb{R})\;\boldsymbol{{\lor}}\;(y_{i}\notin\mathbb{R})\}$?. This is the case when using $D_{S}$ to determine whether the series of digits of $\pi$ is infinitely aperiodic , which in the field of aperiodic, non-rational numbers also called normal (without, in this case, having any relation to the Gauss pdf) [Sevcik2016, Sevcik2017b] :
@@
For any digit $c$ (on the scale $g$) and each natural number $n\in\mathbb{N}$, we denote $I(c,n)$ the number of those digits rm the sequence $c_{1},c_{2},\ldots,c_{n}$, which are equal to $c$. Then
$\lim_{n\to\infty}\left[\dfrac{I(c,n)}{n}\right]=\dfrac{1}{g}$
For each of the $9$ values $\neq 0$ of $c$, then the number $x$ called normal on the scale of $g$. A number that is natural on the $g$ scale is called absolutely natural [22]. For a number with base $10$, the definition implies that $c_{i}\in\mathbb{N}[0,9]$ if it must be true for any number $x\in\mathbb{R}$ as $N\rightarrow\infty$. Therefore, various authors use statistical tests such as comparing the frequencies of each digit in the sequences of the decimals of $\pi$, and the frequencies of various combinations of decimal digits [19]. This approach is unsatisfactory since all possible combinations cannot be evaluated, and is limited by the $N$ values of other series studied, as well as by the size of sub samples of which $N$ was selected to perform statistical tests.
The fractal analysis described [Sevcik2016, Sevcik2017b] considered the set of decimals of $\pi$, and calculated its approximate fractal dimension $D_{S}$ for $N=10^{1},10^{2},10^{3},\ldots,10^{9}$ and showed that